# Diversity of disc viscosities can explain the period ratios of resonant and non-resonant systems of hot super-Earths and mini-Neptunes

### v1

@@ Appendix E Planetary masses
The planetary mass determines not only the migration speed, but are also important for the stability of planetary systems (e.g. Gladman 1993; Chambers et al. 1996). In Fig. 13 we show the cumulative mass distribution of our planets after 100 Myr of evolution. Clearly, the masses of our low viscosity planetary systems are too low compared to the inferred masses from the Kepler observations (using the mass-radius relationship of Wolfgang et al. (2016)). The high viscosity simulation of Izidoro et al. (2021), on the other hand, match very well with the observations, especially for the masses of the chains.
In our model, the only difference between the high viscosity and low viscosity simulations is the migration velocity. The viscosity enters also into the pebble isolation mass (Bitsch et al. 2018b), but is set to the same value in all our simulations to allow an easier comparison. Thus, the difference in mass only arises from the difference in migration velocity. As shown in Fig. 8, the planets formed in low viscosity environment migrate slower and can not reach the inner edge of the disc towards the end of the disc’s lifetime, if they start to migrate from exterior to AU (see also Brasser et al. 2017). However, in this region, the pebble isolation mass starts to increase dramatically (see right in Fig. 13). Therefore the majority of the planets that form via pebble accretion in the low viscosity environments that migrate to the inner disc only reach up to 3 Earth masses. As pebble accretion is very efficient, nearly all planets reach this mass, explaining the lower mass tail in the left of Fig. 13. The larger planetary masses are caused by collisions after the gas disc phase during the instability phase of the systems.
In contrast, the planets formed in simulations with higher viscosity can migrate from much further out into the inner disc region due to their faster migration velocity (Fig. 8). This result in a steady influx of planets that have reached higher masses due to the higher pebble isolation mass exterior to a 6-7 AU. Furthermore, these planets are massive enough that they can collide already during the gas disc-phase (Izidoro et al. 2021), which is not the norm for the planets formed in the low viscosity simulations. Consequently, the planets formed in the high viscosity simulations in the inner disc feature higher planetary masses, even though the formula for the pebble isolation mass is the same.
@@ Appendix F Period ratios
We show in Fig. 14 a mixture of 2% stable chains and 98% systems that underwent instabilities. This mixture was used in the past high viscosity simulations of Izidoro et al. (2021), as it gave the best fit to the observations (indicated by the purple line). The low viscosity simulations still fail to reproduce the observed period-ratios, as their period ratios remain too tight.
Appendix G Resonant chains of 2 planet systems
In order to investigate the influence of viscosity on the period ratio of planetary systems, we conduct a simple experiment. We take systems with 2 equal mass planets and let them migrate through a discs, where only the viscosity varies between the different set-ups. Different viscosities can cause different migration velocities (see appendix A) due to the change of the strength of the entropy driven corotation torque (e.g. Paardekooper et al. 2011; Bitsch et al. 2013). In particular, high viscosity environments can lead to outward migration, if the disc’s surface density and temperature gradients are in the right regime. In fact, high viscosities will lead to outward migration in our used disc model (Bitsch et al. 2015a), which complicates our simple experiment. We thus use a simple power law disc model for this experiment:
@@ Appendix G Resonant chains of 2 planet systems
The here observed trend is reflected in our main simulations, where the simulations at low viscosity with low masses show wider final period ratios compared to the simulations with higher masses and higher viscosities. We note though that the exact period ratio in a multi-planet system is more complicated as in the 2 planetary case, due to the increase in eccentricities caused by the gravitational interaction of multiple planets that can influence the exact resonance configuration.

### v2

@@ Appendix E Planetary masses
The planetary mass determines not only the migration speed, but are also important for the stability of planetary systems (e.g. Gladman 1993; Chambers et al. 1996). In Fig. 13 we show the cumulative mass distribution of our planets after 100 Myr of evolution. Clearly, the masses of our low-viscosity planetary systems are too low compared to the inferred masses from the Kepler observations (using the mass-radius relationship of Wolfgang et al. (2016)). The high-viscosity simulation of Izidoro et al. (2021), on the other hand, match very well with the observations, especially for the masses of the chains.
In our model, the only difference between the high viscosity and low-viscosity simulations is the migration velocity. The viscosity enters also into the pebble isolation mass (Bitsch et al. 2018b), but is set to the same value in all our simulations to allow an easier comparison. Thus, the difference in mass only arises from the difference in migration velocity. As shown in Fig. 8, the planets formed in low-viscosity environment migrate slower and can not reach the inner edge of the disc towards the end of the disc’s lifetime, if they start to migrate from exterior to AU (see also Brasser et al. 2017). However, in this region, the pebble isolation mass starts to increase dramatically (see right in Fig. 13). Therefore the majority of the planets that form via pebble accretion in the low-viscosity environments that migrate to the inner disc only reach up to 3 Earth masses. As pebble accretion is very efficient, nearly all planets reach this mass, explaining the lower mass tail in the left of Fig. 13. The larger planetary masses are caused by collisions after the gas disc phase during the instability phase of the systems.
In contrast, the planets formed in simulations with higher viscosity can migrate from much further out into the inner disc region due to their faster migration velocity (Fig. 8). This result in a steady influx of planets that have reached higher masses due to the higher pebble isolation mass exterior to a 6-7 AU. Furthermore, these planets are massive enough that they can collide already during the gas disc-phase (Izidoro et al. 2021), which is not the norm for the planets formed in the low-viscosity simulations. Consequently, the planets formed in the high-viscosity simulations in the inner disc feature higher planetary masses, even though the formula for the pebble isolation mass is the same.
@@ Appendix F Period ratios
We show in Fig. 14 a mixture of 2% stable chains and 98% systems that underwent instabilities. This mixture was used in the past high-viscosity simulations of Izidoro et al. (2021), as it gave the best fit to the observations (indicated by the purple line). The low-viscosity simulations still fail to reproduce the observed period ratios, as their period ratios remain too tight.
Appendix G Resonant chains of two-planet systems
In order to investigate the influence of viscosity on the period ratio of planetary systems, we conduct a simple experiment. We take systems with 2 equal mass planets and let them migrate through a discs, where only the viscosity varies between the different set-ups. Different viscosities can cause different migration velocities (see Appendix A) due to the change of the strength of the entropy driven corotation torque (e.g. Paardekooper et al. 2011; Bitsch et al. 2013). In particular, high-viscosity environments can lead to outward migration, if the disc’s surface density and temperature gradients are in the right regime. In fact, high viscosities will lead to outward migration in our used disc model (Bitsch et al. 2015a), which complicates our simple experiment. We thus use a simple power law disc model for this experiment:
@@ Appendix G Resonant chains of two-planet systems
The here observed trend is reflected in our main simulations, where the simulations at low viscosity with low masses show wider final period ratios compared to the simulations with higher masses and higher viscosities. We note though that the exact period ratio in a multi-planet system is more complicated as in the two-planet case, due to the increase in eccentricities caused by the gravitational interaction of multiple planets that can influence the exact resonance configuration.